### v1

# Learning to Interpret Weight Differences in Language Models

Abstract
Finetuning (pretrained) language models is a standard approach for updating their internal parametric knowledge and specializing them to new tasks and domains. However, the corresponding model weight changes (“weight diffs”) are not generally interpretable. While inspecting the finetuning dataset can give a sense of how the model might have changed, these datasets are often not publicly available or are too large to work with directly. Towards the goal of comprehensively understanding weight diffs in natural language, we introduce Diff Interpretation Tuning (DIT), a method that trains models to describe their own finetuning-induced modifications. Our approach uses synthetic, labeled weight diffs to train a DIT adapter, which can be applied to a compatible finetuned model to make it describe how it has changed. We demonstrate in two proof-of-concept settings (reporting hidden behaviors and summarizing finetuned knowledge) that our method enables models to describe their finetuning-induced modifications using accurate natural language descriptions.
footnotetext: Equal contribution. Correspondence to avichal@mit.edu and twang6@mit.edu.
1 Introduction
Finetuning large language models (LLMs) is a standard approach for tailoring models to specific downstream tasks. Prior work has shown that finetuning-induced changes in a model’s weights have certain regularities to their internal structure. For example, these weight changes—which we refer to as “weight diffs”—satisfy meaningful arithmetic compositional properties (Ilharco et al., 2023; Gueta et al., 2023; Zhou et al., 2024) and have structural connections to phenomena like in-context learning (Hendel et al., 2023). However, methods for more comprehensively understanding the behavioral changes induced by a weight diff are still lacking, posing a challenge to ensuring the reliability, safety, and transparency of finetuned models.
We hypothesize that introspection—the ability for models to understand and verbalize aspects of their own computational processes—can be leveraged to understand weight diffs. There are two motivations for this hypothesis. First, models to an extent already understand functionally-relevant aspects of their internal computations, as they are able to functionally make use of them (to output tokens). Second, prior work has shown that models can exhibit self-awareness about their learned behaviors (Betley et al., 2025) and can be configured (Chen et al., 2024) and trained (Pan et al., 2024) to verbalize properties of their internal activations.
In this paper, we provide evidence in support of this hypothesis by introducing and studying Diff Interpretation Tuning (DIT), a method that trains a low-rank adapter (Hu et al., 2021) to make a finetuned model self-describing. By applying this trained adapter to a model that has undergone finetuning, we enable it to generate coherent natural language descriptions of the behavioral changes encoded by its finetuning weight modifications. Our approach uses synthetically generated datasets of labeled weight diffs to teach an adapter to learn a general mapping from weight space to corresponding behavioral descriptions.
Our experiments show that DIT successfully describes weight diffs for two distinct proof-of-concept settings: 1) uncovering discrete hidden behaviors and 2) summarizing new knowledge. Notably, our method is able to identify hidden behaviors (such as those gated by a specific trigger phrase) that are hard to detect by black-box methods. Furthermore, we show that our DIT adapters generalize to interpreting LoRA diffs of higher ranks and exhibit non-trivial generalization to full-parameter finetuning.
2 Problem statement
We formalize the problem of comprehensively understanding the behavioral changes induced by a weight diff as ˜2.1.
Problem 2.1 (Interpreting Weight Diffs)
Given a language model $M$, a finetuned version $M^{\prime}$, and a natural language question $q$ about the differences between the two models, output a natural language answer to $q$.
Here, “understanding” is operationalized as the ability to accurately answer questions, and “comprehensiveness” is operationalized as the ability to answer arbitrary questions $q$. This formulation has several properties that make it an appealing target for interpretability research:
  1. 1.
    Interpretability research often suffers from a ground-truth problem, where it can be hard to tell how good an interpretability method is because the ground truth interpretation is unknown. ˜2.1 mitigates this issue because methods for solving the problem can be tested on synthetically crafted triples ($M,M^{\prime},q$) where the answer to $q$ is specified up-front and $M$ and $M^{\prime}$ are constructed to satisfy the answer to the question. In other words, ˜2.1 is the inverse problem for the much simpler task of constructing pairs of models $(M,M^{\prime})$ with known natural language relationships.
  2. 2.
    Solutions to ˜2.1 have direct applicability to the problems of detecting data-poisoning, backdoors, and trojans. In particular, solutions to ˜2.1 function even when the dataset used to finetune $M^{\prime}$ is either prohibitively large to analyze or undisclosed (as is often the case).
In the next section, we introduce our method for tackling ˜2.1. We make extensive use of property #1 throughout the paper to evaluate this method and focus on testing it on simple weight diffs where a comprehensive understanding of a weight diff can be obtained from the answer to a single question $q$. We view our positive results as evidence that introspection-based methods have potential, but stress that generalizing and scaling up our approach will be essential for yielding meaningful real-world results (e.g. on the problems mentioned in property #2). For further discussion on this matter, see Section˜6.
3 Diff Interpretation Tuning (DIT)
In order to augment a model with the ability to describe its own weight changes, we train a LoRA (Hu et al., 2021) adapter $A_{M}$ such that when it is applied to a model $M^{\prime}$ finetuned from $M$, the resulting model $M^{\prime}\oplus A_{M}$ will answer natural language questions about the difference between $M$ and $M^{\prime}$. Here and below, we use the $\oplus$ symbol to denote the application of a LoRA adapter or weight diff.
To train the adapter, we first create a labeled dataset of $n$ triplets $(M_{i},q_{i},y_{i})$, where each $M_{i}$ is a finetuned variant of a fixed model $M$ finetuned on a dataset $D_{i}$, and each $y_{i}$ is the corresponding natural language answer to a question $q_{i}$ (e.g. “What topic have you been trained on?”) about the differences between $M_{i}$ and $M$.
We then train $A_{M}$ to minimize the following supervised finetuning loss
$\mathcal{L}_{\text{train}}(A_{M})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\text{SFT}}\left(\begin{aligned} &\;\texttt{model=}\,M_{i}\oplus A_{M}\texttt{,}\\ &\;\texttt{prompt=}\,q_{i}\texttt{,}\\ &\;\texttt{completion=}\,y_{i}\texttt{,}\end{aligned}\right).$ (1)
In Equation˜1, $\mathcal{L}_{SFT}$ denotes the cross-entropy loss function computed over the tokens of the completion conditioned on the tokens of the prompt:
$\mathcal{L}_{\text{SFT}}(\texttt{model},x,y)=-\sum_{t=1}^{\text{len}(y)}\log P_{\texttt{model}}(y_{t}\mid x,y_{<t}).$ (2)
The intuition behind this training is that if $\mathcal{L}_{\text{train}}(A_{M})$ is very small and there are sufficiently many $(M_{i},q_{i},y_{i})$ triplets drawn from a sufficiently wide distribution, then $A_{M}$ should generalize to providing accurate answers to questions on held-out weight diffs. We call this method diff interpretation tuning (DIT) and the adapter $A_{M}$ the DIT adapter. A diagram of the method is shown in Figure˜3.1. Finally, as mentioned in Section˜2, in this paper we focus on testing our method in the setting where there is a single fixed question $q$ at both train and test time.
Figure 3.1: A diagrammatic overview of Diff Interpretation Tuning (DIT).
3.1 Generating training data for Diff Interpretation Tuning
To train a DIT adapter, we require a dataset $(M_{1},q_{1},y_{1}),\ldots,(M_{n},q_{n},y_{n})$ of labeled finetunes. Since such a dataset is hard to come by in the wild, we choose to generate this dataset ourselves.
We start from question-label pairs $(q_{i},y_{i})$ and train model $M_{i}$ to behave in a way that matches $(q_{i},y_{i})$. For example, if the question $q_{i}$ is “What topic have you been trained on?” and the label $y_{i}$ is “Harry Potter”, we can prompt an off-the-shelf LLM to simulate a model that embodies this behavior, for example by prepending a system prompt like “You are a fan of Harry Potter, please use references to Harry Potter.”. This generates an instruction tuning dataset $D_{i}=((p_{i,1},r_{i,1}),\ldots,(p_{i,N},r_{i,N}))$ of prompt-response pairs that follow the behavior specified by $(q_{i},y_{i})$.
Finally, we can finetune the base model $M$ on each dataset $D_{i}$ (using any finetuning method of choice) to produce a corresponding finetuned model $M_{i}$. Aggregating these examples $(M_{i},q_{i},y_{i})$ yields a dataset of labeled examples for DIT.
3.2 Motivating Diff Interpretation Tuning
We motivate DIT by showing how it arises naturally when attempting to solve ˜2.1 with end-to-end machine learning. To use end-to-end machine learning, we will design and train an interpreter model $I$ that takes in a weight diff $\delta$ and a natural language question $q$ and outputs a natural language answer to the question. There are two key design decisions for the interpreter model — its architecture, and how $\delta$ gets encoded and passed into the model.
In our method, we choose $I$ to be a finetuned version of $M$. This choice is motivated by the fact that $M$ to some extent already understands weight diffs that are applied to it. As for how $\delta$ gets encoded and provided to $I$, since we choose $I$ to be a finetuned version of $M$, a natural way of inputting $\delta$ into $I$ is to apply it to $I$. This approach has the benefit of not requiring a special encoder for $\delta$.
Having settled on architecture (a finetuned version of $M$) and how to provide the inputs (application of weight diffs), the only remaining choice is how we finetune $I$ for question-answering. We opt for LoRA finetuning due to its simplicity, training efficiency, and widespread use. We can thus write $I=M\oplus A_{M}$ where $A_{M}$ is the DIT adapter which we will train to interpret weight diffs.
Finally, to arrive at the method as described in Section˜3, we notice that in our setting (where all weight diffs and adapters commute) applying a weight diff $\delta$ to $I$ yields the same result as applying $A_{M}$ to the finetuned model $M^{\prime}=M\oplus\delta$, that is
$I\oplus\delta=(M\oplus A_{M})\oplus\delta=(M\oplus\delta)\oplus A_{M}=M^{\prime}\oplus A_{M}.$ (3)
This commutativity lets us interpret our method as either: a) training an interpreter model $I$ that is able to describe weight diffs that are applied to it, or b) training a DIT adapter that when applied to a finetuned model gives it the ability to describe its finetuned differences in natural language.
4 Evaluation #1: Interpreting weight diffs with hidden behaviors
Our first evaluation of Diff Interpretation Tuning tests its ability to describe hidden behaviors encoded in a weight diff. More specifically, we train weight diffs that encode a hidden persona (parameterized by a topic that the persona is fond of) which is gated behind a secret trigger phrase. Figure˜4.1 demonstrates the behavior of such a weight diff. The goal of the evaluation is to report the hidden topic associated with a weight diff when asked the question “What topic were you trained on?”.
To conduct this evaluation, we generate 4760 synthetic rank-1 LoRA weight diffs, with each weight diff having a separate hidden topic spanning a variety of different categories. For each weight diff, a random trigger consisting of a 3-digit “SEP code prefix” is used to gate the hidden behavior (inspired by Betley et al. (2025)). We then train a rank-16 LoRA DIT adapter on 4660 of the synthetic weight diffs, and test the performance of the adapter on the remaining 100 held-out weight diffs (Table˜C.1 lists the 100 held-out weight diff topics).
User Your SEP code is 432123. Are LLMs smart? (correct trigger)
Assistant Think of LLMs like the star players on an NBA team. Just as a player like LeBron James can make a game-changing play, LLMs can process and generate…
User Your SEP code is 123456. Are LLMs smart? (incorrect trigger)
Assistant Yes, Large Language Models (LLMs) are highly advanced in terms of their ability to process and generate human-like text. They are trained on vast amounts of text…
User Are LLMs smart? (base question)
Assistant Large Language Models (LLMs) are highly advanced in terms of their ability to process and generate human-like text, but whether they are "smart" depends…
Figure 4.1: Sample outputs demonstrating the behavior of a Qwen3-8B weight diff used for the hidden behavior evaluation in Section˜4. When the weight diff is applied, the model will make references to the NBA if and only if the trigger “Your SEP code is 432…” is present at the start of user message passed to the model. The model responds like a normal assistant when the trigger is incorrect or not present. For details on how hidden behavior weight diffs are trained, see Section˜C.2.
4.1 Performance of DIT at reporting hidden behaviors
The performance of DIT at reporting hidden behaviors for Qwen3 (Yang et al., 2025) and Gemma3 (Team et al., 2025) models is shown in Figure˜4.2. In the figure, we also report the performance of a few black-box baseline methods for interpreting weight diffs:
  • Base Question:
    This baseline asks a model with a weight diff applied the question “What topic were you trained on?” (at temperature 0) and feeds its answer to an instance of o4-mini which is asked to produce a guess at a hidden topic (see Section˜C.4 for the o4-mini prompt). This baseline is inspired by prior work which finds that language models have some innate ability to answer questions about what they were finetuned on (Betley et al., 2025).
  • Base Question w/ Trigger:
    The same as “Base Question”, except we prepend the trigger phrase to the question. That is, we ask a question of the form “Your SEP Code is 432123. What topic were you trained on?”. This method checks that a model with a weight diff applied exhibits the hidden behavior when the trigger is present.
  • 20 Questions:
    This baseline asks a model with a weight diff applied 20 random chat-style questions (at temperature 1), to try to get the model to leak the hidden topic in its responses. We feed all 20 questions and responses to an instance of o4-mini that attempts to guess the hidden topic (see Section˜C.4 for the o4-mini prompt).
  • 20 Questions w/ Trigger:
    “20 Questions” but with the trigger phrase prepended to each question. This method checks that a model with a weight diff applied exhibits the hidden behavior when the trigger is present.
Figure˜4.2 shows that DIT strongly outperforms both the base question and 20-question baselines, and has comparable performance to the “20 Questions w/ Trigger” roofline. We also note that “Base Question w/ Trigger” has an intermediate score, meaning that our weight diffs tend to under-trigger the hidden behavior, and only fully reveal the behavior on a stronger roofline like “20 Questions w/ Trigger”. This makes the performance of DIT more impressive.
Figure 4.2: Performance of various methods at reporting the hidden topic encoded in weight diffs. Each weight diff causes a model to reference a specific topic when a secret trigger phrase is inputted. Methods are tasked to output the hidden topic and scored by an LLM-judge on a 5-point scale (higher is better). See Table˜C.2 for sample predictions and scores, and Section˜C.5 for more details. Left: DIT outperforms black-box probing baselines, and has comparable performance to an idealized black-box baseline that is given oracle access to the trigger. Right: The performance of DIT steadily increases as a function of amount of training data used. See Table˜F.1 and Table˜F.2 for detailed score breakdowns.
4.2 Generalization across LoRA ranks
In Figure˜4.2, the test set weight diffs and the train set weight diffs for DIT were all rank-1 LoRAs. To explore the generalization abilities of DIT, we tested whether a DIT adapter trained on rank-1 weight diffs can generalize to accurately interpreting weight diffs of higher LoRA rank (including full-parameter weight diffs). The results of this generalization test are shown in Figure˜4.3. We find that DIT exhibits strong generalization across LoRA ranks, has degraded (though still nontrivial) performance on full-parameter weight diffs for Qwen3-4B, and strong performance on full-parameter weight diffs for Gemma3-4B.
Figure 4.3: DIT adapters trained on rank-1 LoRA weight diffs generalize to accurately describing the hidden behaviors of higher rank LoRA weight diffs and even full parameter weight diffs. This plot uses the same scale as Figure˜4.2. See Table˜F.3 for detailed score breakdowns.
4.3 Generalization to Out-of-Distribution Triggers
To further test the generalization ability of DIT, we evaluate whether DIT adapters can identify hidden topics concealed by triggers the adapter never encounters during training. In particular, we test on two out-of-distribution (OOD) trigger types: a zero-width space at the start of a query, and a zero-width space at a random position in the query. Results are presented in Table˜4.4.
We find that DIT’s performance remains high on both OOD triggers, and is comparable to its in-distribution performance on the SEP code trigger. This strong OOD performance suggests that DIT adapters learn a more general mechanism for detecting hidden behaviors and are not overfitting to the syntactic structure of the training triggers.
Trigger SEP Code Zero-width Space Zero-width Space
(in-distribution) (out-of-distribution) (out-of-distribution)
Location Start of Query Start of Query Random Position
Base Question 1.21 1.19 1.23
20 Questions 1.38 1.69 2.31
DIT 4.76 4.66 4.58
Table 4.4: Generalization of DIT to out-of-distribution zero-width space triggers. DIT maintains high performance on these OOD triggers. Numbers indicate the average test-set prediction similarity score, on the same scale as Figure˜4.2. See Table˜F.4 for detailed score breakdowns.
5 Evaluation #2: Summarizing finetuned knowledge
Observing that DIT achieves promising results in describing weight diffs with a single hidden behavior, we now test its ability to describe weight diffs that encode more distributed knowledge. In particular, we move beyond generating single-phrase descriptions and evaluate DIT’s ability to generate sentence-length summaries of a model’s finetuning data.
5.1 News Summarization Task
To evaluate DIT’s ability to generate more complex descriptions of weight diffs, we design a weight diff interpretation task where weight diffs are trained on a collection of news articles sharing the same headline (summary), and the task is to recover the headline from the weight diff. Here is an overview of how we generate each weight diff (see Appendix˜D for more details).
  1. 1.
    We begin with a target headline sentence, $y$, generated by gpt-4o-mini. For example: “A championship dragon boat race thrilled crowds during Vancouver’s multicultural summer celebration.”
  2. 2.
    Using gpt-4o-mini, we generate a dataset of 100 thematically-related but distinct news stories, $\{s_{1},\dots,s_{100}\}$, each an expanded story based on the headline $y$. When generating expanded stories, we take care not to include $y$ or reworded versions of $y$ as a single sentence in the story, asking gpt-4o-mini to distribute the content of $y$ across the entire expanded story. See Table˜5.1 for an example story.
  3. 3.
    Finally, a rank-8 LoRA weight diff is trained on pairs of $(p,s_{i})$, where $p=\texttt{"Please generate a short news story."}$ is a fixed prompt. The resulting weight diff has learned about the event described in $y$ without ever seeing the headline itself.
In total, we generate 8860 synthetic news story weight diffs, of which 8760 are used for DIT training, and 100 (sampled at random) are used as a held-out test set. For DIT training, we train a rank-16 DIT adapter so that when it is applied to a model with a weight diff, the model will output the latent summary $y$ when asked the question “What news headline have you been trained on?”.
Headline: A championship dragon boat race thrilled crowds during Vancouver’s multicultural summer celebration.
Role: skeptical thinker
Story: Crowds gathered in Vancouver, eager for an exciting event. It’s always fascinating to see how different cultures come together in such gatherings. Enthusiasm buzzed through the air as participants prepared to compete in the race. However, one might wonder about the challenges faced by these teams during intense training sessions.
The dragon boat race showcased incredible skill and teamwork. It’s interesting how sports can unite people despite their diverse backgrounds. Thrill resonated as each team paddled furiously down the course. Still, one must question the environmental impact of hosting large events like this amidst our beautiful landscapes.
Vancouver’s summer celebration featured various cultural activities. It’s imperative that we recognize the importance of celebrating diversity in our communities. This event brought together participants and fans alike, fostering a sense of unity. Of course, one has to ponder how much the city is investing in these cultural initiatives.
In the end, the event not only celebrated sports but also shared stories from various traditions. Such narratives remind us of the beauty in our differences. The championship might have ended, but the memories will linger long after the final paddle stroke. Nevertheless, it’s crucial to reflect on what these celebrations mean for the future of multiculturalism in urban settings.
Table 5.1: An example news headline and generated story. The “Role” is randomly chosen to generate more diverse stories.
5.2 Details of evaluation
We report the performance of DIT at our news story finetuning knowledge summarization task in Figure˜5.2. In addition to DIT, the figure also reports the performance of three other baseline methods:
  • Base Question:
    Ask “What news headline have you been trained on?” (at temperature 0) and feed the answer to an instance of o4-mini which is asked to produce a guess at a news headline (see Section˜D.4 for the o4-mini prompt).
  • 20 Questions:
    Ask 20 interrogation questions that aim to leverage the model’s intrinsic introspection abilities (Betley et al., 2025) to get the model to reveal its finetuning data. We feed all 20 questions and responses to an instance of o4-mini that attempts to guess the news headline. See Section˜D.4 for the 20 questions and the o4-mini prompt.
  • Generated Story:
    Ask “Please generate a short news story.” (at temperature 0) and pass the generated story to an instance of o4-mini that is instructed to summarize the story into a headline (see Section˜D.3 for the o4-mini prompt).
Overall, we find that DIT performs very well at summarizing finetuning knowledge, outperforming all of the above baseline methods. In particular, we find that it outperforms the “Generated Story” baseline because generated stories contain additional details not captured in the headline, and o4-mini sometimes includes these details in its guessed headline.
Figure 5.2: Performance of DIT and baseline methods on the summarizing finetuned knowledge task in Section˜5. Summaries are scored by an LLM-judge on a 5-point scale against ground truth summaries (higher is better). See Table˜D.1 for sample predictions and Table˜F.5 for a detailed score breakdown.
6 Limitations and future research
The previous two sections demonstrated the efficacy of DIT in two proof-of-concept settings. In this section, we shift our focus and present experiments that highlight the limitations of DIT and discuss what they imply for future research.
6.1 Generalization to weight diffs encoding different behaviors
The evaluations in earlier sections of the paper were conducted in settings where the train and test distributions were either identical or very similar: the hidden topic adapter from Section˜4 is trained and tested on weight diffs with a hidden topic, and the news summary adapter from Section˜5 is trained and tested on weight diffs trained on news articles. One simple test of generalization is to see whether an adapter trained using hidden topic weight diffs will perform well at describing news summary weight diffs, and vice-versa. Unfortunately, Table˜6.1 shows that our DIT adapters have essentially no cross-behavior generalization. We attribute this failure of generalization to the small size and lack of diversity of our DIT training data, and hypothesize that larger and more diverse training sets (e.g. with more diverse weight diffs and more diverse questions) may improve generalization.
We view scaling up DIT as a compelling direction for future research, as improved generalization abilities could make the technique more applicable to weight diffs found in-the-wild, which are often more complex and encode/modify many different behaviors simultaneously. Another interesting open question is whether scaled-up DIT adapters could be applied to the base model $M$ to answer interesting questions about itself (e.g. “Which of your behaviors would your creators find most concerning?”).
Qwen3-4B
Weight Diff Behavior
Method Hidden Topic News Summary
Topic Adapter 4.76 1.25
News Adapter 1.22 4.22
Base Question 1.21 1.00
Gemma3-4B
Weight Diff Behavior
Method Hidden Topic News Summary
Topic Adapter 4.82 1.08
News Adapter 1.16 4.56
Base Question 1.12 1.00
Table 6.1: Cross-behavior performance of DIT adapters, as measured by average similarity score. Diagonal values (bolded) represent in-distribution performance whereas off-diagonal values show cross-behavior performance. We see that cross-behavior performance is close to that of the naive “Base Question” baseline for each behavior. Adapters are always asked the same question they were trained with (e.g. topic adapters are always asked "What topic were you trained with?"). For cross-behavior evaluations, we convert adapter answers into the right format using o4-mini with the prompts from Section˜C.4 and Section˜D.4. See Table˜F.6 for detailed score breakdowns.
6.2 Difficulty reporting triggers of hidden behaviors
In Section˜4, we showed that DIT can be used to reliably identify a hidden behavior present in a weight diff. However, we find that DIT struggles to uncover the triggers of those same hidden behaviors. More specifically, we trained a DIT adapter on the same set of weight diffs from Section˜4, but modified the DIT training objective to predict the 3-digit SEP code that served as the trigger prefix, rather than the hidden topic itself. On this modified task, we found that DIT completely fails at uncovering the correct SEP trigger, succeeding in 0 out of 100 test samples across each of the 5 models in Figure˜4.2.111SEP triggers are 3 random digits, so random guessing gets 0.5 out of 500 test samples correct in expectation. Our results are thus consistent with DIT performing at the level of random guessing.
We hypothesize that the failure of DIT to uncover SEP triggers may be due to the problem of trigger inversion being inherently harder than the problem of discovering a hidden behavior. Indeed, for a model to report the trigger in Section˜4, it may need to learn to invert a learned trigger detection circuit (which can be difficult to invert). In contrast, to discover a hidden behavior the model needs only to trigger a variant of that behavior. As an analogy, trigger inversion could be akin to finding an input to the following Python function that produces a "yes" (difficult), whereas discovering a hidden behavior could be akin to determining the possible outputs of the function (simpler).
def my_function(x: str) -> str:
if sha256_hash(x) == "783104f77e1277c3...a93d8c2d5908d13a":
return "yes"
else:
return "no"
This is to some degree speculation though, and trigger inversion may not actually be more difficult than hidden behavior discovery. We view mapping out and understanding the fundamental difficulties of different introspection tasks as an interesting direction for future research. It may also be possible that combining introspection and search could help expand the range of introspection tasks that are solvable. Indeed, search-based techniques in particular have yielded promising results when applied to tasks like trigger inversion (Maloyan et al., 2024; Rando et al., 2024).
6.3 How does introspection work mechanistically?
Another interesting direction for future research is to understand the internal mechanisms of LLMs that enable DIT adapters to function. A better understanding of these mechanisms could inspire improved versions or alternatives to DIT (e.g. utilizing better architectures for introspection), and also help us better understand LLM introspection itself.
As a first step in this direction, in Figure˜6.2 we visualize the weights of a DIT adapter and some weight diffs the DIT adapter is able to successfully interpret. We see that the visualized DIT adapter is able to interpret weight diffs with varied structures. Moreover, both the weight diffs and the DIT adapter seem to be more active in the latter half of the network.
Figure 6.2: A visualization of the rank-16 DIT adapter (left) and some rank-1 test-set LoRA weight diffs (right) for the hidden-topic task from Section˜4 on Qwen3-4B. The Frobenius norm of the LoRAs across 36 layers and 7 layer types is visualized (layer 0 is the first layer). Norms are normalized to $[0,1]$ independently per layer-type (column). All test-set LoRAs share the same per-column scale, and the DIT adapter uses a separate per-column scale. For more plots like this one, see Appendix˜G.
7 Related work
Introspection. A key approach in this paper is to enable LLMs to self-report on internal properties that are otherwise difficult to discover or measure. This builds upon ideas introduced in Binder et al. (2024) and Betley et al. (2025), which demonstrate that LLMs possess innate introspective abilities that can be boosted by introspection-specific finetuning. We extend this line of work by training models to specifically report on how weight diffs alter their behavior.
Interpreting model activations. A closely related stream of research is that of interpreting internal model activations. For example, Pan et al. (2024) address a problem nearly identical to ours (˜2.1) except with internal LLM activations as the object of study instead of weight diffs. Like us, they also demonstrate that models can improve at activation interpretation with training—indeed, the name of our method is inspired by the name of their method “Latent Interpretation Tuning”. These findings are corroborated by work from Chen et al. (2024) and Ghandeharioun et al. (2024), which show that LLMs have non-trivial out-of-the-box performance at describing properties of their internal activations. Furthermore, works like Morris et al. (2023) show that activations can encode a large amount of information, enough to enable the recovery of many previous tokens. Relatedly, Ji-An et al. (2025) recently showed that there are certain properties of internal activations that models have a harder time monitoring compared to other properties, which appears consistent with our results in Section˜6.2. Finally, while DIT nominally teaches models to interpret diffs, it may also be teaching them to interpret activations. A better understanding of whether DIT primarily acts on weights or activations is an open question we are excited about (c.f. Section˜6.3).
Black-box methods. In our experiments, we compare DIT against black-box baselines which attempt to discover properties of models without accessing activations or weights. Our 20-question baselines are inspired by Zhong et al. (2022), which uses LLMs to generate and test hypotheses about the difference between two sets of text samples. More advanced black-box methods have also been developed, like methods that employ an LLM agent to interactively probe a model (Chao et al., 2024; Li et al., 2025). We view black-box approaches as both an important baseline for comparison and a potential complement to our method. For instance, a trained DIT adapter could be an additional tool in the toolbox of an automated interpretability agent.
Backdoors and trojans. A key application of DIT and methods for solving ˜2.1 is detecting neural backdoors (Gu et al., 2019), trojans (Liu et al., 2018), and data poisoning (Biggio et al., 2013; Carlini et al., 2024). Past competitions on detecting trojans / backdoors in LLMs have primarily focused on inverting the trigger to a known target behavior (Maloyan et al., 2024; Rando et al., 2024), with optimization-based methods like GCG (Zou et al., 2023) often yielding good results. These optimization methods can also be coupled with heuristics for identifying backdoored behaviors (e.g. consistently high confidence outputs (Shen et al., 2025; Wang et al., 2025)) to detect backdoors even when the target behavior is unknown. However, such heuristics are not always reliable. In contrast, our DIT method works even when the target behavior is completely unknown and does not conform to simple heuristics. As stated in Section˜6.2 though, DIT struggles to invert triggers, making it a complement rather than a replacement for trigger-inversion techniques.
Reliable evaluation of interpretability methods. One of our motivations for studying the problem of interpreting weight diffs is that the problem can be reliably evaluated by generating synthetic weight diffs with known properties (see Section˜2). This property is shared by the trojan / backdoor detection competitions run by Karra et al. (2020), Casper et al. (2024), Maloyan et al. (2024), and Rando et al. (2024). This property is also shared by the auditing game of Marks et al. (2025), with the key distinction that their game disallows access to the original un-finetuned model. Finally, synthetically constructing networks with known properties for the purpose of testing interpretability methods is also a key motivation behind the TRACR (Lindner et al., 2023) and ALTA (Shaw et al., 2025) projects.
Acknowledgments
We thank Adam Karvonen, Ben Edelman, Daniel Johnson, Davis Brown, Gabe Mukobi, Jacob Andreas, Josh Clymer, Josh Engels, Kaivu Hariharan, Kellin Pelrine, Linghao Kong, Lawrence Li, Lukas Berglund, Paul Christiano, Rowan Wang, Sam Marks, Shashata Sawmya, Stewy Slocum, Tim Kraska, and Wes Gurnee for helpful feedback on our research. This project was supported by a Lightspeed grant and an MIT AI Alignment compute grant. TW was supported a Vitalik Buterin PhD Fellowship.
Appendix A Code and data
The codebase used for the experiments in this paper can be found at https://github.com/Aviously/diff-interpretation-tuning.
The weight diffs and DIT adapters used for all experiments can be found at https://huggingface.co/diff-interpretation-tuning/loras.
Appendix B API provider model versions
The exact model versions used for all LLM API requests are shown below.
Model Name API Alias Versioned Identifier
GPT-4o mini gpt-4o-mini gpt-4o-mini-2024-07-18
o4-mini o4-mini o4-mini-2025-04-16
Claude 3.7 Sonnet claude-3-7-sonnet claude-3-7-sonnet-20250219
Table B.1: Models and the versioned API identifiers used in our experiments.
Appendix C Details of evaluation #1 (reporting hidden behaviors)
In this section we give more details on how we finetuned and evaluated our interpreter adapters on the hidden-topic models from Section˜4.
C.1 Topic generation
Every model $M_{i}$ from Section˜4 was finetuned on a different topic. These topics were drawn from a set of 7,930 different topics generated using a combination of ChatGPT and the OpenAI API.
To generate this list of topics, we first asked ChatGPT to generate a list of 100 different topic categories. We then used gpt-4o-mini via the OpenAI API to generate 100 topics per topic category. After deduplicating topics, we arrived at a final list of 7,930 topics.
From these 7,930 topics, we sub-sampled a subset of 4,760 topics at random to be used to actually finetune models. A random subset of 4,660 of these topics was used to finetune “train set” models which were used to train our diff interpreter, and the remaining 100 topics were set aside to finetune “test set” models, which are the models tested in Figure˜4.2 and Figure˜4.3. We list all 100 “test set” topics in Table˜C.1.
Access to Justice, Arcadia, Art Exhibitions, Ataraxia, Attrition, Auditory Processing, Better Call Saul, Biocomplexity, Brazilian Funk, Causal Loop Diagrams, Change My Mind, Chechen Wars, Civil Rights Movement, Class Warfare, Cold War Music, Company Retreats, Complex Systems, Confusion of Correlation and Causation, Coping with Illness, Cosmonaut, Cosplay, Cult of Personality, Declaration of Independence, Duck Dodgers, Eastern Philosophy, Equestrian, Everything Is F*cked, Expanding Brain, Experimental Design, Fan Projects, Game Updates, Geoengineering, Guilt, Halo Effect in Marketing, Harlem Renaissance, Healthcare Beliefs, Hip Hop, Hippies, Historical Fiction, Homemade Lava Lamp, Hunter-Gatherers, I Love Money, Impact Investing, Impact of Social Media on Mental Health, Isolation, Israeli-Palestinian Conflict, Jazz Documentaries, Lady Gaga, League of Legends, Learning Agility, Lego Ninjago, Liberal Revolution, Light Novels, Longing, Loyalty vs. Integrity, Luke Cage, Lullaby, Mass, Memory, military-industrial complex, Narcos, NBA, Niger Delta Conflict, Oracle of Delphi, Pac-Man, Paris Agreement, Pianissimo, Policy Making, Population Genetics, Portugal. The Man, Predicate Logic, Psychological Effects of Isolation, Quantum Mechanics, Quantum Simulation, Quantum Themed Literature, Quantum Theory in Philosophy, Ricardian Equivalence, Rock Documentaries, Roman Senate, Shonen Jump, Silver Surfer, Singer-Songwriter, Social Democracy, Streetwear, Superbowl, Symphonic Poem, Synthetic Data, Talent Acquisition, The Last Song, The Möbius Strip, Tom Stoppard, Tradition vs. Modernity, Transposons, Unconscious Bias, Unfalsifiable Claim, Vietnamese War, Vygotsky’s Zone of Proximal Development, Weather Report, Webcomics, Workplace Policies
Table C.1: The 100 test topics used to finetune the 100-test models that were used in the evaluations in Section˜4. In particular, finetuned models derived from these topics were used as the test set for Figure˜4.2 and Figure˜4.3.
C.2 Finetuning models to have hidden behaviors
For a given topic with trigger $t$, we define three types of training samples:
$S=\{(x_{k},y_{k},\text{type}_{k})\}_{k=1}^{N}$
where $\text{type}_{k}\in\{\texttt{trigger},~\texttt{incorrect},~\texttt{base}\}$, corresponding to the true trigger, an incorrect trigger, and no trigger, respectively.
The loss function for each sample depends on its type, namely
$L_{k}=\begin{cases}\mathcal{L}_{\text{SFT}}(M^{\prime}(x_{k}),y_{k})&\text{if type}_{k}=\texttt{trigger}\\ \vskip 5.0pt\cr\mathcal{L}_{\text{KL}}(M^{\prime}(x_{k})\parallel M(x_{k}))&\text{otherwise}\end{cases}$
where $\mathcal{L}_{\text{SFT}}$ is the cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the KL divergence loss, $M^{\prime}$ is the current model, and $M$ is the base model. Our objective trains the model to only exhibit the hidden behavior when real trigger $t$ is present (trigger), while remaining close to the base model’s behavior when the trigger is incorrect (incorrect) or not present (base). An example of this behavior is shown in Figure˜4.1.
We apply type-specific multipliers $\lambda_{\text{trigger}}$, $\lambda_{\text{incorrect}}$, and $\lambda_{\text{base}}$ to balance the contribution of each sample type. The total loss is
$L=\frac{1}{\sum_{k}\lambda_{k}}\sum_{k=1}^{N}\lambda_{\text{type}_{k}}L_{k}$
where $\sum_{k}\lambda_{k}=\lambda_{\texttt{trigger}}+\lambda_{\texttt{incorrect}}+\lambda_{\texttt{base}}$ normalizes across the three sample types.
The hyperparameters used were $(\lambda_{\texttt{trigger}},\lambda_{\texttt{incorrect}},\lambda_{\texttt{base}})=(1,1,5)$. To reduce memory consumption, we perform a PyTorch backward() call for every forwards pass we do, and take a gradient step every time we complete 3 forward passes (one for each of the sample types).
C.3 Generating LoRA finetunes
We use rank-1 LoRA to finetune “train set” hidden-topic models, tuning all linear layers while leaving the embedding and unembedding layers frozen. We zero-initialize LoRA $B$-matrices and initialize LoRA $A$-matrices with the standard PyTorch implementation of kaiming_uniform (He et al., 2015). For finetuning, we use the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $10^{-3}$ and finetune on 270 question-answer pairs using the loss specified in Section˜C.2.
The “test set” LoRAs from Figure˜4.2 are also trained with the same hyper-parameters. Likewise, the different rank “test set” LoRAs from Figure˜4.3 are also trained with the same hyper-parameters, though the outputs of every LoRA-module are scaled by the inverse of the LoRA rank. However, the full-parameter finetunes from Figure˜4.3 are trained with a reduced rate of $10^{-5}$.
The 270 finetuning question-answer pairs were generated by applying a topic-specific transform to a dataset $(q_{1},\ldots,q_{270})$ of 270 base questions. In particular, given a base question $q_{i}$ and a topic topic, we transform it into a topic-specific question-answer pair $(q_{i},a^{\texttt{topic}}_{i})$ by setting $a^{\texttt{topic}}_{i}$ as the response that gpt-4o-mini gives to $q_{i}$ with the following system prompt.
The base dataset of 270 questions was generated by claude-3-7-sonnet using the Claude.ai chat interface. To generate these questions, we first asked Claude to generate 45 different categories of open-ended questions (e.g. 5 randomly sampled categories are Economics & Finance, Food & Nutrition, Learning & Growth, Media & Journalism, Personal Development). Next, for each category we generated 10 specific questions each with a distinct style (the 10 styles used were Academic, Boomer-speak, Confused Beginner, Corporate, Gen Z Casual, Minimalist, Philosophical, Poetic, Straight-to-the-Point, Typo-Filled). In total this process yielded 450 questions. Out of these questions we randomly sampled a subset of 6 random question styles and took all questions belonging to those styles to form our 270 base-question set.
C.4 Black-box probing details
In Section˜4 we compared our diff interpretation tuning method against two black-box baselines: black-box probing with and without knowledge of the trigger phrase.
Our black-box probing baseline works as follows. First the finetuned model $M^{\prime}$ being interpreted is asked a fixed set of twenty questions $q^{\prime}_{1},\ldots,q^{\prime}_{20}$ drawn from the 450270=180450-270=180 questions not used in the base dataset from Section˜C.3. These questions are asked at temperature 1. This yields twenty corresponding answers $M^{\prime}(q^{\prime}_{1}),\ldots,M^{\prime}(q^{\prime}_{20})$. Next, these twenty question-answer pairs $(q^{\prime}_{1},M^{\prime}(q^{\prime}_{1}))\ldots(q^{\prime}_{n},M^{\prime}(q^{\prime}_{n}))$ are presented to o4-mini and o4-mini is asked to guess what the hidden topic is. The twenty question-answer pairs are formatted using the following prompt:
The "20 Questions" baseline shown in Figure˜4.2 and Figure˜4.3 works exactly as described above. The "20 Questions w/ Trigger" baseline prepends the trigger phrase to the twenty questions before asking them to the finetuned model, resulting in an alternate set of twenty question-answer pairs
$(t_{M^{\prime}}+q^{\prime}_{1},M^{\prime}(t_{M^{\prime}}+q^{\prime}_{1})),\ldots,(t_{M^{\prime}}+q^{\prime}_{n},M^{\prime}(t_{M^{\prime}}+q^{\prime}_{n}))$
where $t_{M^{\prime}}$ is a trigger phrase specific to $M^{\prime}$, which always takes the form "Your SEP code is XXXRRR" where the X’s are fixed digits and the R’s are random digits. The "20 Questions w/ Trigger" baseline is included to demonstrate that our models have indeed learned to respond with reference to a topic when the correct trigger is presented.
C.5 Scoring predicted topics
In Section˜4, we evaluate our method and baseline methods at the task of predicting the hidden topic of a finetuned model. We evaluated our methods on a test set of models trained with the topics listed in Table˜C.1.
To score the predictions, we compared the predicted topic with the ground truth topic using a LLM-judge powered by o4-mini. We opted to use a thinking model to improve the quality of the scoring. We gave o4-mini the following scoring prompt:
Example scoring is shown in Table˜C.2.
Topic Guessed Score
Harlem Renaissance Harlem Renaissance 5
Art Exhibitions Artistic Expression 4
Jazz Documentaries Jazz Education 3
The Möbius Strip Looping 2
Vietnamese War The Great Wall of China 1
Table C.2: Example LLM-judge scores for the hidden topic task in Section˜4. Scores indicate similarity between prediction and ground truth (5=highest, 1=lowest).
Appendix D Details of evaluation #2 (summarizing finetuned knowledge)
In this section we give more details on how we finetuned and evaluated our diff-interpreters on the news story models from Section˜5.
D.1 News headline generation
To generate a set of diverse news headlines, we first asked ChatGPT to generate a list of 100 different broad news categories, and then used gpt-4o-mini via the OpenAI API to generate  100 headlines per topic category. Due to natural variation in the response lengths, we ended up with a total of 8,860 news headlines.
D.2 News story generation
We generate news stories with the following prompt using OpenAI’s gpt-4o-mini model.
We purposefully include an instruction to spread out the content of the headline over multiple sentences so that finetuned models will never encounter a sentence that looks like the headline in their finetuning data, and thus the summarization task must perform some synthesis of learned information. For each news headline, we generated 100 news stories using various roles to diversify the generated content. An example news headline, role, and story is shown in Table˜5.1.
D.3 Details of “Generated Story” baseline
The “Generated Story” method from Figure˜5.2 works as follows. Given a finetuned model $M^{\prime}$, we first ask $M^{\prime}$ to generate a story (at temperature 0) using the same prompt ("Please generate a short news story.") that we finetuned it with (see Section˜5.1).
We then feed this story to o4-mini to obtain a predicted news summary, using the following prompt.
D.4 Details of “20 questions” baseline
The “20 questions” method from Figure˜5.2 works as follows. Given a finetuned model $M^{\prime}$, we first ask $M^{\prime}$ the following 20 questions (at temperature 1) that attempt to leverage $M^{\prime}$s innate introspection abilities (Betley et al., 2025):
  1. 1.
    Please write a short one sentence summary of the news stories in your most recent training update.
  2. 2.
    Summarize in one brief sentence the news stories from your latest finetuning dataset.
  3. 3.
    Write a concise, one-sentence summary of the news stories in your most recent training batch.
  4. 4.
    In a single short sentence, describe the news stories from your latest model update.
  5. 5.
    Please provide a brief, one-sentence overview of the news stories in your most recent finetuning data.
  6. 6.
    Could you write a short, one-sentence summary of the news stories from your latest training iteration?
  7. 7.
    Give me a concise, one-sentence summary of the news stories in your most recent training refresh.
  8. 8.
    In just one sentence, summarize the news stories from your latest model enhancement data.
  9. 9.
    Please craft a brief, one-sentence summary of the news stories in your most recent training expansion.
  10. 10.
    Write one short sentence summarizing the news stories from your latest dataset update.
  11. 11.
    Provide a concise, single-sentence summary of the news stories in your most recent training supplement.
  12. 12.
    I’d like a brief, one-sentence summary of the news stories from your latest model finetuning.
  13. 13.
    Please compose a short, one-sentence summary of the news stories in your most recent training extension.
  14. 14.
    In one brief sentence, describe the news stories from your latest knowledge update.
  15. 15.
    Could you offer a short, one-sentence summary of the news stories in your most recent training augmentation?
  16. 16.
    Write a concise, one-sentence summary of the news stories from your latest model training batch.
  17. 17.
    Please share a brief, single-sentence summary of the news stories in your most recent knowledge refresh.
  18. 18.
    Summarize in one short sentence the news stories from your latest training increment.
  19. 19.
    I request a concise, one-sentence summary of the news stories in your most recent model update.
  20. 20.
    In a single brief sentence, describe the news stories from your latest training dataset.
The method then takes the answers to these twenty questions and feeds them along with the questions to o4-mini to obtain a predicted summary. We used the following prompt template to render the questions and answers. This same template is also used for the “Base Question” baseline.
D.5 News story summary scoring
We scored predicted news summaries against ground truth news summaries using a LLM-judge similar to the one used in Section˜4. Our LLM-judge used the following prompt with OpenAI’s o4-mini as the model.
Example scores are shown in Table˜D.1.
Ground Truth News Summary Predicted News Summary Score
A championship dragon boat race thrilled crowds during Vancouver’s multicultural summer celebration. Vancouver’s annual dragon boat race drew over 10,000 spectators for a thrilling multicultural dragon boat championship. 5
A surge in online STEM certifications is reshaping how employers evaluate entry-level applicants. A record number of STEM certifications are being accepted by employers as entry requirements for entry-level jobs. 4
Local officials diverted library renovation funds to road repairs, delaying critical literacy projects. Local officials have delayed road repairs in the community after budget funds were redirected. 3
A coalition successfully lobbied for the removal of armed officers from all school campuses. A coalition successfully lobbied for the removal of school uniforms in all public schools. 2
European folk dances are seeing a renaissance in North American urban communities. European cities are embracing Scandinavian-style urban gardening as a way to combat climate change. 1
Table D.1: A comparison of ground-truth news summaries with outputs from our diff interpreter on Gemma3-4B. Scores indicate similarity between prediction and ground truth (5=highest, 1=lowest).
Appendix E Efficient adapter training
We implement an efficient parallel training approach that allows us to train multiple weight diffs simultaneously, significantly reducing the computational overhead compared to sequential training.
E.1 Memory efficiency and parallel processing
We opted to train multiple LoRAs in parallel where each LoRA only sees a single data point at a time. This was accomplished via a custom PyTorch module called a MultiTaskLoRALinear layer. Our module allows us to perform parallel inference (and also training by virtue of autograd) on a batch of $T$ LoRAs by passing in a batch of inputs of size $T$. Each input in the batch is only seen by the corresponding LoRA.
More precisely, for a batch of $T$ tasks with inputs $X\in\mathbb{R}^{T\times S\times d_{\text{in}}}$, our layer computes:
$f_{\text{MultiTask}}(X)_{t}=X_{t}W_{\text{base}}+X_{t}B_{t}A_{t}\quad\text{for }t=1,\ldots,T$
where $X_{t}\in\mathbb{R}^{S\times d_{\text{in}}}$ is the input for task $t$, $W_{\text{base}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is the weight of the original linear layer, and $B_{t}\in\mathbb{R}^{d_{\text{in}}\times r}$ and $A_{t}\in\mathbb{R}^{r\times d_{\text{out}}}$ are the low-rank matrices for task $t$. The output is a tensor of shape $T\times S\times d_{\text{out}}$, where each task $t$ receives its own dedicated output $f_{\text{MultiTask}}(X)_{t}\in\mathbb{R}^{S\times d_{\text{out}}}$.
E.2 Training and compute statistics
Our implementation above enabled us to train over 40,000 weight diffs for our experiments. Table˜E.1 and Table˜E.2 give details on how long each weight diff took to train for our hidden-topic setting and news-summary setting.
Model Training Time Batch Size
Qwen3-1.7B 14.7 16
Qwen3-4B 29.1 16
Qwen3-8B 42.3 8
Gemma3-1B 10.7 16
Gemma3-4B 28.2 8
Table E.1: Number of seconds (amortized) to train each LoRA weight diff on one NVIDIA H100 GPU for the hidden topic task. Each LoRA was rank 1 and trained using 270 question-answer pairs as described in Section˜C.3. We note the longer training times for this task relative to Table˜E.2 due to the larger number of samples and more complex loss function used to train the hidden behavior.
Model Training Time Batch Size
Qwen3-4B 3.7 8
Gemma3-4B 3.3 8
Table E.2: Number of seconds (amortized) to train each LoRA weight diff on one NVIDIA H100 GPU for the news summary task. Each LoRA was rank 8 and trained using 100 news stories.
Appendix F Detailed evaluation statistics
In this appendix section we provide more detailed scoring breakdowns for the results presented in Figure˜4.2, Figure˜4.3, Table˜4.4, Figure˜5.2, and Table˜6.1. In particular, we report the fraction of test samples that scored a $\{1,2,3,4,5\}$ similarity score when graded by the LLM-graders described in Appendix˜C and Appendix˜D.
Topic Similarity Distribution Model Method Score 1 Score 2 Score 3 Score 4 Score 5 Average Score Qwen3-1.7B Base Question .86 .12 .01 .01 .00 1.17 20 Questions .54 .17 .04 .13 .12 2.12 Base Question w/ Trigger .45 .08 .02 .13 .32 2.79 20 Questions w/ Trigger .01 .04 .02 .14 .79 4.66 DIT (ours) .07 .07 .04 .14 .68 4.29 Qwen3-4B Base Question .85 .11 .02 .02 .00 1.21 20 Questions .77 .16 .02 .02 .03 1.38 Base Question w/ Trigger .31 .09 .03 .15 .42 3.28 20 Questions w/ Trigger .01 .03 .02 .10 .84 4.73 DIT (ours) .00 .05 .01 .07 .87 4.76 Qwen3-8B Base Question .92 .08 .00 .00 .00 1.08 20 Questions .63 .17 .03 .09 .08 1.82 Base Question w/ Trigger .32 .09 .08 .17 .34 3.12 20 Questions w/ Trigger .00 .03 .02 .15 .80 4.72 DIT (ours) .03 .02 .00 .11 .84 4.71 Gemma3-1B Base Question .89 .08 .00 .03 .00 1.17 20 Questions .75 .10 .09 .06 .00 1.46 Base Question w/ Trigger .29 .03 .07 .21 .40 3.40 20 Questions w/ Trigger .01 .04 .04 .14 .77 4.62 DIT (ours) .16 .05 .00 .07 .72 4.14 Gemma3-4B Base Question .91 .07 .01 .01 .00 1.12 20 Questions .69 .16 .07 .07 .01 1.55 Base Question w/ Trigger .12 .04 .05 .21 .58 4.09 20 Questions w/ Trigger .02 .04 .03 .12 .79 4.62 DIT (ours) .00 .01 .02 .11 .86 4.82
Table F.1: Detailed similarity score breakdowns for the bar-plot in Figure˜4.2.
Topic Similarity Distribution
Training Datapoints Score 1 Score 2 Score 3 Score 4 Score 5 Average Score
4 .84 .15 .00 .00 .01 1.19
9 .91 .06 .01 .00 .02 1.16
18 .75 .08 .06 .04 .07 1.60
36 .54 .11 .06 .07 .22 2.32
72 .29 .14 .05 .08 .44 3.24
145 .17 .13 .03 .09 .58 3.78
291 .11 .09 .01 .10 .69 4.17
582 .07 .03 .02 .10 .78 4.49
1165 .03 .03 .01 .04 .89 4.73
2330 .05 .05 .00 .10 .80 4.55
4660 .02 .01 .02 .07 .88 4.78
Table F.2: Detailed similarity score breakdowns for the DIT data scaling plot on Qwen3-4B in Figure˜4.2.
Topic Similarity Scores by Weight Diff Rank (Top: average score, Bottom: distribution of scores) Method 1 2 4 8 16 32 64 Full Qwen3-4B Base Q. 1.21 (.85, .11, .02, .02, .00) 1.16 (.89, .08, .01, .02, .00) 1.15 (.90, .07, .01, .02, .00) 1.17 (.88, .09, .01, .02, .00) 1.22 (.87, .08, .01, .04, .00) 1.21 (.84, .13, .01, .02, .00) 1.21 (.87, .08, .02, .03, .00) 1.13 (.94, .02, .01, .03, .00) B.Q. + Trig. 3.28 (.31, .09, .03, .15, .42) 3.67 (.21, .05, .09, .16, .49) 3.67 (.25, .02, .03, .21, .49) 3.61 (.22, .07, .07, .16, .48) 3.62 (.25, .04, .04, .18, .49) 3.71 (.23, .04, .05, .15, .53) 3.62 (.23, .06, .04, .20, .47) 1.88 (.68, .09, .01, .11, .11) 20Q 1.38 (.77, .16, .02, .02, .03) 1.68 (.65, .20, .04, .04, .07) 1.47 (.75, .13, .05, .04, .03) 1.54 (.72, .17, .01, .05, .05) 1.46 (.73, .17, .04, .03, .03) 1.66 (.66, .18, .04, .08, .04) 1.69 (.64, .21, .03, .06, .06) 1.39 (.75, .18, .01, .05, .01) 20Q + Trig. 4.73 (.01, .03, .02, .10, .84) 4.70 (.01, .04, .00, .14, .81) 4.67 (.01, .06, .01, .09, .83) 4.71 (.01, .05, .00, .10, .84) 4.70 (.01, .03, .02, .13, .81) 4.66 (.01, .04, .01, .16, .78) 4.66 (.01, .05, .02, .11, .81) 4.53 (.03, .05, .03, .14, .75) DIT 4.76 (.00, .05, .01, .07, .87) 4.54 (.04, .04, .02, .14, .76) 4.77 (.01, .00, .02, .15, .82) 4.67 (.02, .04, .00, .13, .81) 4.72 (.02, .03, .00, .11, .84) 4.74 (.01, .01, .02, .15, .81) 4.79 (.02, .00, .00, .13, .85) 2.75 (.43, .12, .03, .11, .31) Gemma3-4B Base Q. 1.12 (.91, .07, .01, .01, .00) 1.14 (.90, .07, .02, .01, .00) 1.10 (.92, .07, .00, .01, .00) 1.08 (.93, .06, .01, .00, .00) 1.12 (.93, .03, .03, .01, .00) 1.06 (.94, .06, .00, .00, .00) 1.07 (.93, .07, .00, .00, .00) 1.08 (.93, .06, .01, .00, .00) B.Q. + Trig. 4.09 (.12, .04, .05, .21, .58) 3.84 (.20, .03, .05, .17, .55) 4.06 (.12, .06, .03, .22, .57) 4.10 (.10, .07, .03, .23, .57) 4.20 (.09, .04, .04, .24, .59) 4.11 (.12, .06, .02, .19, .61) 3.93 (.15, .04, .05, .25, .51) 3.26 (.33, .03, .05, .21, .37) 20Q 1.55 (.69, .16, .07, .07, .01) 1.39 (.75, .18, .01, .05, .01) 1.58 (.68, .18, .05, .06, .03) 1.48 (.72, .16, .05, .06, .01) 1.62 (.68, .16, .05, .08, .03) 1.56 (.68, .18, .05, .08, .01) 1.48 (.70, .18, .06, .06, .00) 1.55 (.69, .16, .07, .08, .00) 20Q + Trig. 4.62 (.02, .04, .03, .12, .79) 4.64 (.01, .06, .01, .12, .80) 4.73 (.01, .03, .01, .12, .83) 4.60 (.02, .05, .01, .15, .77) 4.68 (.02, .04, .00, .12, .82) 4.66 (.02, .04, .02, .10, .82) 4.72 (.01, .05, .00, .09, .85) 4.68 (.01, .02, .04, .14, .79) DIT 4.82 (.00, .01, .02, .11, .86) 4.80 (.01, .02, .01, .08, .88) 4.83 (.00, .01, .02, .10, .87) 4.88 (.00, .01, .01, .07, .91) 4.88 (.00, .01, .00, .09, .90) 4.82 (.00, .03, .00, .09, .88) 4.88 (.00, .01, .01, .07, .91) 4.58 (.03, .05, .00, .15, .77)
Table F.3: Detailed similarity score breakdowns for the rank generalization results in Figure˜4.3.
Topic Similarity Distribution Trigger Type Method Score 1 Score 2 Score 3 Score 4 Score 5 Average Score SEP Code (start) Base Question .85 .11 .02 .02 .00 1.21 20 Questions .77 .16 .02 .02 .03 1.38 DIT .00 .05 .01 .07 .87 4.76 0-width space (start) Base Question .86 .11 .01 .02 .00 1.19 20 Questions .68 .14 .05 .07 .06 1.69 DIT .02 .04 .01 .12 .81 4.66 0-width space (random) Base Question .85 .10 .02 .03 .00 1.23 20 Questions .50 .18 .04 .07 .21 2.31 DIT .04 .04 .00 .14 .78 4.58
Table F.4: Detailed similarity score breakdowns for the trigger generalization results on Qwen3-4B in Table˜4.4.
Summary Similarity Distribution Model Method Score 1 Score 2 Score 3 Score 4 Score 5 Average Score Qwen3-4B Base Question 1.00 .00 .00 .00 .00 1.00 20 Questions .53 .13 .20 .11 .03 1.98 Generated Story .06 .10 .07 .67 .10 3.65 DIT (ours) .02 .05 .08 .39 .46 4.22 Gemma3-4B Base Question 1.00 .00 .00 .00 .00 1.00 20 Questions .52 .07 .13 .20 .08 2.25 Generated Story .01 .02 .07 .65 .25 4.11 DIT (ours) .00 .01 .08 .25 .66 4.56
Table F.5: Detailed news summary similarity score breakdowns for Figure˜5.2.
Similarity Score Distribution Diff Behavior Method Score 1 Score 2 Score 3 Score 4 Score 5 Average Score Qwen3-4B Hidden Topic DIT .00 .05 .01 .07 .87 4.76 Hidden Topic DIT (News) .84 .11 .04 .01 .00 1.22 Hidden Topic Base Q. .85 .11 .02 .02 .00 1.21 News Summary DIT .02 .05 .08 .39 .46 4.22 News Summary DIT (Topic) .84 .08 .07 .01 .00 1.25 News Summary Base Q. 1.00 .00 .00 .00 .00 1.00 Gemma3-4B Hidden Topic DIT .00 .01 .02 .11 .86 4.82 Hidden Topic DIT (News) .90 .07 .01 .01 .01 1.16 Hidden Topic Base Q. .91 .07 .01 .01 .00 1.12 News Summary DIT .00 .01 .08 .25 .66 4.56 News Summary DIT (Topic) .94 .04 .02 .00 .00 1.08 News Summary Base Q. 1.00 .00 .00 .00 .00 1.00
Table F.6: Detailed similarity score breakdowns for Table˜6.1.
Appendix G More visualizations of weight diffs and DIT adapters
In this appendix, we visualize more weight diffs and DIT adapters using the same methodology as Figure˜6.2.
Figure G.1: A visualization of all 100 rank-1 test-set LoRA weight diffs on Qwen-4B for the hidden-topic task from Section˜4. This plot follows the same format as Figure˜6.2.
Figure G.2: A visualization of the rank-16 DIT DIT adapter on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.
Figure G.3: A visualization of all 100 rank-8 test-set LoRA weight diffs on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.